OpenStudy (anonymous):

Solve for x: |x − 2| + 10 = 12

OpenStudy (anonymous):

|x − 2| + 10 = 12 Transfer 10 to right side, what would you get?

OpenStudy (anonymous):

You will get |x − 2| = 2

OpenStudy (anonymous):

Now \[ x − 2 = \pm2\] Choosing one plus and minus sign you can solve it yourself

OpenStudy (anonymous):

for positive charge we can write x-2+10=12 =>x+8=12 =>x=12-8 =>x=4 for negative charge we can write -(x-2)+10=12 =>-x+2+10=12 =>-x=12-12 =>x=0

OpenStudy (anonymous):

x-2=2 or x-2=-2 i.e. x= 4 or x=0 is the ans